Mattola Company is giving each of its employees a holiday bonus of $900 on December 13,...

  1. Mattola Company is giving each of its employees a holiday bonus of $900 on December 13, 20-- (a nonpayday). The company wants each employee's check to be $900. The supplemental tax percent is used.

    Nobody has capped for OASDI prior to the bonus check.

    a. What will be the gross amount of each bonus if each employee pays a state income tax of 2.8% (besides the other payroll taxes)? You may need to add one penny to the gross so that net bonus exactly equals $900. Round your calculations and final answers to the nearest cent.

    $

    b. What would the net amount of each bonus check be if the company did not gross-up the bonus? Round your intermediary calculations to the nearest cent.

    $

    Feedback

    a. Divide desired net pay by: 1 (representing 100%) minus all tax rates (using decimals) that will be deducted from gross pay. This will give the required gross pay needed to give desired net pay.

    Take the resulting gross pay and multiply by the respective tax rates to prove net pay. Please note, due to rounding, the net bonus can be off by a cent. Add or subtract this cent to gross pay to ensure that once withheld taxes are subtracted, the net check equals the bonus amount. Use 22% as FIT percentage on supplemental earnings.

    b. Gross pay minus deductions = net pay. Use 22% as FIT percentage on supplemental earnings.
Answer #1
a)
Assume the gross bonus to be paid to every employee is 100x
State income tax is 2.8%
So, State income tax = 2.8% of 100x = 2.8x
Remaining bonus after paying state income tax = 100x - 2.8x = 97.2x
Now FIT% is 22%
So, FIT = 22% of 97.2x = 21.384x
So net bonus amount = 97.2x - 21.384x = 75.816x
According to the statement, net bonus to be paid to employees = 900
Keeping both equal, 75.816x = 900
So, gross bonus = 100x = 900 *100/75.816
Gross bonus = $1,187.08
b)
If Gross bonus = $900
State income tax is 2.8%
So, State income tax = 2.8% of 900 = 25.20
Remaining bonus after paying state income tax = 900 - 25.20 = 874.8
Now FIT% is 22%
So, FIT = 22% of 874.8 = 192.456
So net bonus amount = 874.8 - 192.456
Net bonus = $682.34